# SWIFTCODER: Enhancing Code Generation in Large Language Models through Efficiency-Aware Fine-tuning

### v3

@@ 3.3 Dataset Statistics
As shown in Table 1, coding problems in SWIFTCODER have been collected from nine datasets, resulting in a total of 65,710 tasks across five programming languages: Python, C++, Java, Rust, and Go. The dataset encompasses a diverse range of coding challenges, ensuring a comprehensive coverage of various programming concepts and problem-solving techniques. Python has the highest representation in SWIFTCODER, with 33,489 tasks sourced from all nine datasets. This extensive collection of Python tasks allows for effective fine-tuning of LLMs to generate efficient and optimized Python code. C++ and Java also have significant contributions, with 11,547 and 14,726 tasks, respectively. These tasks are primarily sourced from CodeFeed, Glaive, Evol-Ins, Dolphin, and Oss-Ins datasets, providing a robust foundation for fine-tuning LLMs in these popular programming languages. Rust and Go, although having relatively fewer tasks compared to Python, C++, and Java, still have a substantial presence in SWIFTCODER. With 4,270 Rust tasks and 1,678 Go tasks, the dataset enables fine-tuning of LLMs to generate efficient code in these modern and rapidly growing programming languages.
Figure 4 illustrates the efficiency distribution of the dataset for three key metrics: execution time, memory usage, and max memory peak, which compares the distribution of these metrics for both inefficient (canonical solutions provided by the nine datasets) and efficient solutions in the SWIFTCODER. For execution time, the inefficient solutions have a mean value of 1.14s, while the efficient solutions have a significantly lower mean of 0.31s, which indicates that the optimization process has successfully reduced the execution time of the code, resulting in more efficient solutions. Similarly, the memory usage and max memory peak also show a notable difference between inefficient and efficient solutions. For example, inefficient solutions have a mean memory usage of 26.50MBs, while efficient solutions have a much lower mean of 6.03MBs.
The efficiency distribution visualization highlights the effectiveness of the optimization process in creating more efficient solutions across all three metrics. By carefully curating tasks through the multi-step cleaning process and applying SOAP optimization, we have created a dataset that is valuable for training models to generate efficient code. SWIFTCODER provides a diverse range of optimized coding problems, enabling researchers and practitioners to advance the field of code optimization using LLMs.
@@ 4 Experiment
Datasets and Models We evaluate the efficiency and correctness of LLM-generated code on five code generation benchmarks, i.e., EffiBench (Huang et al., 2024b), EvalPlus (HumanEvalPlus and MBPPPlus) (Liu et al., 2024), DS-1000 (Lai et al., 2023), EvoEval (Xia et al., 2024), and HumanEval-X (Zheng et al., 2023). We finetune eight open-source LLMs with SWIFTCODER, including CodeLlama-7b-bf, DeepSeek-Coder-6.7B base and instruct model (DeepSeekAI, 2023), Qwen2.5-Code-7B base and instruct model (Hui et al., 2024), and Qwen2.5-Coder (1.5B, 3B, and 14B).
Fine-tuning Setup We use Llama-factory (Zheng et al., 2024) to fully fine-tune all LLMs with the same setup and train the models using SWIFTCODER. The maximum sequence length is set to 2048 tokens. We use a batch size of 128 and set the learning rate to 5e-6 with a cosine learning rate scheduler and a warmup ratio of 0.03. We fine-tune all LLMs for 4 epochs under the bf16 data type.
Table 2: Code efficiency and pass@1 of LLMs trained with SWIFTCODER on EffiBench using greedy decoding. The percentage in the brackets indicates the extent of the reduction for each respective item. Overlap means the percentage of correct tasks addressed by both SWIFTCODER finetuned LLM and original LLM in total tasks of the dataset. We provide a case example in LABEL:fig:example to demonstrate how SWIFTCODER fine-tuned LLM improves the efficiency of LLM-generated code.
@@
+ SFT 0.15 (58.3%) 1.14 (58.2%) 32.41 (0.0%) 1.00 (0.0%) 6.80 (46.0%) 1.23 (52.1%) 50.1 63.6
Table 3: Code efficiency and pass@1 of LLMs trained with SWIFTCODER on HumanEval-EvalPlus.
@@ 4.1 Evaluation of Python Code
To comprehensively demonstrate the efficiency and correctness of automatically generated code by LLMs with the SFT of SWIFTCODER, we first provide the evaluation of the LLMs in generating Python code, where LLMs are asked to create Python code based on natural language or function signatures with docstrings.
EffiBench is a benchmark used to measure the efficiency and correctness of LLM-generated code in LeetCode tasks. To ensure that the efficiency of LLM-generated code can be measured, the authors construct 100 tests for each task to provide enough testing time. As shown in Table 2, we can observe that for all LLMs, the efficiency of the LLM-generated code has been improved after fine-tuning with SWIFTCODER. For example, the average execution time (ET) for the correct code generated by Qwen2.5-Coder-7B-Instruct and its SWIFTCODER fine-tuned version decreases from 0.31 (s) to 0.16 (s), a reduction of 48.4%. Similarly, the total memory usage (TMU) of LLM-generated code also shows significant decreases. For instance, the TMU of DeepSeek-Coder-6.7B-Instruct decreases from 30.05 (Mb*s) to 9.48 (Mb*s), a larger reduction than the decrease in average execution time for the correct code generated by the same model, which only reduces by 35.3% from 0.34 (s) to 0.22 (s). This indicates that during code generation, both the execution time and memory usage of the SWIFTCODER fine-tuned-LLM-generated code have been improved compared to the code generated by the original models. Furthermore, the memory usage/peak (MU) of DeepSeek-Coder-6.7B-Instruct generated code decreases from 47.26 (Mb) to 36.31 (Mb), a reduction of 23.2%, ensuring that the LLM-generated code can be deployed in memory-constrained scenarios such as embedded systems or edge devices.
Interestingly, we observe that compared to the MU, ET is more widely optimized across all models. This suggests that SWIFTCODER fine-tuning has a more significant impact on reducing the execution time of the generated code than on reducing its memory footprint. Nevertheless, the improvements in execution time and memory usage demonstrate the effectiveness of SWIFTCODER in enhancing the efficiency of LLM-generated code.
Table 4: Code efficiency and pass@1 of Qwen2.5-Coder-7B-Instruct and deepseek-coder-6.7b-instruct fine-tuned using SFT with the SWIFTCODER for the EvoEval dataset.
@@
+ SFT 0.26 (38.1%) 0.99 (42.1%) 61.75 (0.2%) 1.00 (0.0%) 10.86 (31.7%) 0.98 (39.5%) 27.0 41.0
Table 5: Code efficiency and pass@1 of Qwen2.5-Coder-7B-Instruct and deepseek-coder-6.7b-instruct fine-tuned using SFT with the SWIFTCODER for the DS-1000 dataset.
@@
+ SFT 1.2587 (4.5%) 1.0082 (3.8%) 441.7553 (1.9%) 1.0013 (2.0%) 6.8022 (5.6%) 1.0132 (5.6%) 29.10 37.50
Table 6: Code efficiency and pass@1 of Qwen2.5-Coder-7B-Instruct and deepseek-coder-6.7b-instruct fine-tuned using SFT with the SWIFTCODER for the HumanEval-X (CPP and Java) dataset.
@@
+ SFT 0.0048 (36.5%) 0.1231 (35.9%) 4.0142 (50.5%) 0.0908 (49.2%) 0.001900 (30.4%) 0.118972 (33.9%) 6.71 57.93
Table 7: Efficiency comparison of different methods on the HumanEvalPlus dataset. We use the fine-tuned CodeLlama-7b-hf by PIE and Mercury as the baselines to measure the improvement of SWIFTCODER fine-tuned version.
@@
PIE 0.30 (23.1%) 1.47 (24.2%) 61.39 (0.5%) 1.00 (0.0%) 11.28 (11.7%) 1.68 (8.2%) 9.8 19.5
Ours 0.21 (46.2%) 1.03 (46.9%) 61.33 (0.6%) 1.00 (0.0%) 7.17 (43.9%) 1.04 (43.2%) 9.8 31.1
Mercury 0.31 (20.5%) 1.51 (22.2%) 61.94 (-0.4%) 1.00 (0.0%) 10.24 (19.9%) 1.47 (19.7%) 4.3 9.1
Ours 0.21 (46.2%) 1.01 (47.9%) 61.73 (-0.1%) 1.00 (0.0%) 6.95 (45.6%) 0.98 (46.4%) 4.3 31.1
@@
HumanEvalPlus and MBPPPlus As shown in Table 3, we observe that almost all LLMs achieve better efficiency and higher correctness after being fine-tuned with SWIFTCODER. Take HumanEvalPlus as an example, the average execution time (ET) for correct code generated by Qwen2.5-Coder-7B-Instruct and its fine-tuned version decreases from 0.52 (s) to 0.32 (s), a reduction of 38.5%. These improvements demonstrate the effectiveness of SWIFTCODER in optimizing the efficiency of LLM-generated code. Moreover, the pass@1 of Qwen2.5-Coder-7B increases from 40.2% to 78.7%, an improvement of 38.5% after fine-tuning with SWIFTCODER. This indicates that the fine-tuned models not only generate more efficient code but also produce correct code more frequently. Similar to the results in HumanEvalPlus, the efficiency and correctness of SWIFTCODER fine-tuned LLMs also improve in the MBPPPlus dataset. For instance, the average execution time for correct code generated by deepseek-coder-6.7b-base decreases by 36.7%, and the total memory usage (TMU) decreases by 40.5% after fine-tuning.
@@
EvoEval includes 828 programming problems created by prompting GPT-4 to evolve original HumanEval tasks across 5 semantic-altering and 2 semantic-preserving benchmarks, each of which has 100 problems. We conduct experiments on the Tool_Use, Subtle, and Creative benchmarks to evaluate the performance of Qwen2.5-Coder-7B-Instruct and deepseek-coder-6.7b-instruct fine-tuned with SWIFTCODER. As shown in Table 4, both models demonstrate significant improvements in code efficiency after fine-tuning with SWIFTCODER. For the Tool_Use benchmark, the average execution time (ET) for correct code generated by Qwen2.5-Coder-7B-Instruct decreases by 42.9%, and the total memory usage (TMU) decreases by 41.6% after fine-tuning. Similarly, deepseek-coder-6.7b-instruct achieves a 41.5% reduction in ET and a 15.9% reduction in TMU. In the Subtle benchmark, after fine-tuning, Qwen2.5-Coder-7B-Instruct and deepseek-coder-6.7b-instruct exhibit 17.5% and 33.3% reductions in ET, respectively, after fine-tuning. The normalized total memory usage (NTMU) also decreases by 19.9% and 34.5% for the two models. Moreover, the pass@1 rate improves significantly, with Qwen2.5-Coder-7B-Instruct increasing from 55.0% to 72.0% and deepseek-coder-6.7b-instruct increasing from 56.0% to 69.0%. For the Creative benchmark, Qwen2.5-coder-7B-instruct and deepseek-coder-6.7b-instruct achieve 27.5% and 38.1% reductions in ET, respectively, after fine-tuning. The NTMU also decreases by 23.8% and 39.5% for the two models. The pass@1 rate remains the same for Qwen2.5-Coder-7B-Instruct at 44.0% but improves from 31.0% to 41.0% for deepseek-coder-6.7b-instruct.
@@ 4.2 Data Science Programming
DS-1000 is a data science benchmark consisting of 1000 realistic challenges across 7 popular Python data science libraries. We evaluate the efficiency and pass@1 of LLM-generated code for the DS-1000 tasks using Qwen2.5-Coder-7B-Instruct and deepseek-coder-6.7b-instruct, both in their original form and fine-tuned with SWIFTCODER. As shown in Table 5, fine-tuning with SWIFTCODER leads to modest improvements in code efficiency for both models. For Qwen2.5-Coder-7B-Instruct, the average execution time (ET) decreases by 0.8%, and the total memory usage (TMU) decreases by 0.9% after fine-tuning, while the pass@1 rate improves significantly from 17.00% to 35.70%. For deepseek-coder-6.7b-instruct, fine-tuning results in a 4.5% reduction in ET and a 5.6% reduction in TMU, with the memory usage (MU) and normalized memory usage (NMU) also decreasing by 1.9% and 2.0%, respectively, although the pass@1 rate remains the same at 37.50%. While the improvements in code efficiency for the DS-1000 benchmark are less pronounced compared to other benchmarks, the results still demonstrate that fine-tuning with SWIFTCODER can enhance the efficiency of LLM-generated code for data science tasks.
@@ 4.3 Different Programming Language
In addition to evaluating the efficiency and pass@1 of LLM-generated code for Python tasks, we also conduct experiments on the HumanEval-X dataset, where we focus on the C++ and Java subsets, to measure the performance of SWIFTCODER fine-tuned LLMs in a different-language setting. As shown in Table 6, both Qwen2.5-Coder-7B-Instruct and deepseek-coder-6.7b-instruct demonstrate significant improvements in code efficiency and pass@1 rates after fine-tuning with SWIFTCODER for C++ tasks. Qwen2.5-Coder-7B-Instruct achieves a 37.1% reduction in average execution time (ET), a 42.6% reduction in memory usage (MU), and a 53.8% reduction in total memory usage (TMU), with the pass@1 rate improving from 7.32% to 48.17%. Similarly, deepseek-coder-6.7b-instruct exhibits a 45.2% reduction in ET, a 50.0% reduction in MU, and a 40.0% reduction in TMU, with the pass@1 rate increasing from 14.63% to 40.24%. For Java tasks, deepseek-coder-6.7b-instruct achieves a 36.5% reduction in ET, a 50.5% reduction in MU, and a 30.4% reduction in TMU after fine-tuning, with the pass@1 rate improving from 14.63% to 57.93%, matching the performance of the fine-tuned Qwen2.5-Coder-7B-Instruct. These results demonstrate that fine-tuning with SWIFTCODER can significantly enhance the efficiency and correctness of LLM-generated code across multiple programming languages.

### v4

@@ 3.3 Dataset Statistics
As shown in Table 1, coding problems in EffiInstruct have been collected from nine datasets, resulting in a total of 65,710 tasks across five programming languages: Python, C++, Java, Rust, and Go. The dataset encompasses a diverse range of coding challenges, ensuring a comprehensive coverage of various programming concepts and problem-solving techniques. Python has the highest representation in EffiInstruct, with 33,489 tasks sourced from all nine datasets. This extensive collection of Python tasks allows for effective fine-tuning of LLMs to generate efficient and optimized Python code. C++ and Java also have significant contributions, with 11,547 and 14,726 tasks, respectively. These tasks are primarily sourced from CodeFeed, Glaive, Evol-Ins, Dolphin, and Oss-Ins datasets, providing a robust foundation for fine-tuning LLMs in these popular programming languages. Rust and Go, although having relatively fewer tasks compared to Python, C++, and Java, still have a substantial presence in EffiInstruct. With 4,270 Rust tasks and 1,678 Go tasks, the dataset enables fine-tuning of LLMs to generate efficient code in these modern and rapidly growing programming languages.
Figure 4 illustrates the efficiency distribution of the dataset for three key metrics: execution time, memory usage, and max memory peak, which compares the distribution of these metrics for both inefficient (canonical solutions provided by the nine datasets) and efficient solutions in the EffiInstruct. For execution time, the inefficient solutions have a mean value of 1.14s, while the efficient solutions have a significantly lower mean of 0.31s, which indicates that the optimization process has successfully reduced the execution time of the code, resulting in more efficient solutions. Similarly, the memory usage and max memory peak also show a notable difference between inefficient and efficient solutions. For example, inefficient solutions have a mean memory usage of 26.50 MBs, while efficient solutions have a much lower mean of 6.03 MBs.
The efficiency distribution visualization highlights the effectiveness of the optimization process in creating more efficient solutions across all three metrics. By carefully curating tasks through the multi-step cleaning process and applying SOAP optimization, we have created a dataset that is valuable for training models to generate efficient code. EffiInstruct provides a diverse range of optimized coding problems, enabling researchers and practitioners to advance the field of code optimization using LLMs.
@@ 4 Experiment
Datasets and Models We evaluate the efficiency and correctness of LLM-generated code on five code generation benchmarks, i.e., EffiBench (Huang et al., 2024b), EvalPlus (HumanEvalPlus and MBPPPlus) (Liu et al., 2024a), DS-1000 (Lai et al., 2023), EvoEval (Xia et al., 2024), and HumanEval-X (Zheng et al., 2023). We finetune eight open-source LLMs with EffiInstruct, including CodeLlama-7b-bf, DeepSeek-Coder-6.7B base and instruct model (DeepSeekAI, 2023), Qwen2.5-Coder-7B base and instruct model (Hui et al., 2024), and Qwen2.5-Coder (1.5B, 3B, and 14B).
Fine-tuning Setup We use Llama-factory (Zheng et al., 2024) to fine-tune LLMs with fully supervised fine-tuning with the same setup and train the models using EffiInstruct. The maximum sequence length is set to 2048 tokens. We use a batch size of 128 and set the learning rate to 5e-6 with a cosine learning rate scheduler and a warmup ratio of 0.03. We fine-tune all LLMs for 4 epochs under the bf16 data type.
Table 2: Code efficiency and pass@1 of LLMs trained with EffiInstruct on EffiBench using greedy decoding. The percentage in the brackets indicates the extent of the reduction for each respective item. Overlap means the percentage of correct tasks addressed by both EffiInstruct finetuned LLM and original LLM in total tasks of the dataset.
@@
+ EffiInstruct 0.15 (58.3%) 1.14 (58.2%) 32.41 (0.0%) 1.00 (0.0%) 6.80 (46.0%) 1.23 (52.1%) 50.1 63.6
Table 3: Code efficiency and pass@1 of LLMs trained with EffiInstruct on HumanEvalPlus and MBPPPlus.
@@ 4.1 Evaluation of Python Code
To comprehensively demonstrate the efficiency and correctness of automatically generated code by LLMs with the SFT of EffiInstruct, we first provide the evaluation of the LLMs in generating Python code, where LLMs are asked to create Python code based on natural language or function signatures with docstrings.
EffiBench is a benchmark used to measure the efficiency and correctness of LLM-generated code in LeetCode tasks. To ensure that the efficiency of LLM-generated code can be measured, the authors construct 100 tests for each task to provide enough testing time. As shown in Table 2, we can observe that for all LLMs, the efficiency of the LLM-generated code has been improved after fine-tuning with EffiInstruct. For example, the average execution time (ET) for the correct code generated by Qwen2.5-Coder-7B-Instruct and its EffiInstruct fine-tuned version decreases from 0.31 (s) to 0.16 (s), a reduction of 48.4%. Similarly, the total memory usage (TMU) of LLM-generated code also shows significant decreases. For instance, the TMU of DeepSeek-Coder-6.7B-Instruct decreases from 30.05 (Mb*s) to 9.48 (Mb*s), a larger reduction than the decrease in average execution time for the correct code generated by the same model, which only reduces by 35.3% from 0.34 (s) to 0.22 (s). This indicates that during code generation, both the execution time and memory usage of the fine-tuned-LLM-generated code have been improved compared to the code generated by the original models. Furthermore, the memory usage/peak (MU) of DeepSeek-Coder-6.7B-Instruct generated code decreases from 47.26 (Mb) to 36.31 (Mb), a reduction of 23.2%, ensuring that the LLM-generated code can be deployed in memory-constrained scenarios such as embedded systems or edge devices.
Interestingly, we observe that compared to the MU, ET is more widely optimized across all models. This suggests that EffiInstruct fine-tuning has a more significant impact on reducing the execution time of the generated code than on reducing its memory footprint. Nevertheless, the improvements in execution time and memory usage demonstrate the effectiveness of EffiInstruct in enhancing the efficiency of LLM-generated code.
Table 4: Code efficiency and pass@1 of Qwen2.5-Coder-7B-Instruct and deepseek-coder-6.7b-instruct fine-tuned using SFT with the EffiInstruct for the EvoEval dataset.
@@
+ EffiInstruct 0.26 (38.1%) 0.99 (42.1%) 61.75 (0.2%) 1.00 (0.0%) 10.86 (31.7%) 0.98 (39.5%) 27.0 41.0
Table 5: Code efficiency and pass@1 of Qwen2.5-Coder-7B-Instruct and deepseek-coder-6.7b-instruct fine-tuned using SFT with the EffiInstruct for the DS-1000 dataset.
@@
+ EffiInstruct 1.2587 (4.5%) 1.0082 (3.8%) 441.7553 (1.9%) 1.0013 (2.0%) 6.8022 (5.6%) 1.0132 (5.6%) 29.10 37.50
Table 6: Code efficiency and pass@1 of Qwen2.5-Coder-7B-Instruct and deepseek-coder-6.7b-instruct fine-tuned using SFT with the EffiInstruct for the HumanEval-X (CPP and Java) dataset.
@@
+ EffiInstruct 0.0048 (36.5%) 0.1231 (35.9%) 4.0142 (50.5%) 0.0908 (49.2%) 0.001900 (30.4%) 0.118972 (33.9%) 6.71 57.93
Table 7: Efficiency comparison of different methods on the HumanEvalPlus dataset. We use the fine-tuned CodeLlama-7b-hf by PIE and Mercury as the baselines to measure the improvement of EffiInstruct fine-tuned version.
@@
CodeLlama-7b-hf 0.39 1.94 61.68 1.00 12.78 1.83 9.8 12.2
+ PIE (All) 0.30 (23.1%) 1.47 (24.2%) 61.39 (0.5%) 1.00 (0.0%) 11.28 (11.7%) 1.68 (8.2%) 9.8 19.5
+ EffiInstruct 0.21 (46.2%) 1.03 (46.9%) 61.33 (0.6%) 1.00 (0.0%) 7.17 (43.9%) 1.04 (43.2%) 9.8 31.1
CodeLlama-7b-hf 0.39 1.94 61.69 1.00 12.78 1.83 4.3 12.2
+ Mercury 0.31 (20.5%) 1.51 (22.2%) 61.94 (-0.4%) 1.00 (0.0%) 10.24 (19.9%) 1.47 (19.7%) 4.3 9.1
+ EffiInstruct 0.21 (46.2%) 1.01 (47.9%) 61.73 (-0.1%) 1.00 (0.0%) 6.95 (45.6%) 0.98 (46.4%) 4.3 31.1
@@
HumanEvalPlus and MBPPPlus As shown in Table 3, we observe that almost all LLMs achieve better efficiency and higher correctness after being fine-tuned with . Take HumanEvalPlus as an example, the average execution time (ET) for correct code generated by Qwen2.5-Coder-7B-Instruct and its fine-tuned version decreases from 0.52 (s) to 0.32 (s), a reduction of 38.5%. These improvements demonstrate the effectiveness of in optimizing the efficiency of LLM-generated code. Moreover, the pass@1 of Qwen2.5-Coder-7B increases from 40.2% to 78.7%, an improvement of 38.5% after fine-tuning with . This indicates that the fine-tuned models not only generate more efficient code but also produce correct code more frequently. Similar to the results in HumanEvalPlus, the efficiency and correctness of fine-tuned LLMs also improve in the MBPPPlus dataset. For instance, the average execution time for correct code generated by deepseek-coder-6.7b-base decreases by 36.7%, and the total memory usage (TMU) decreases by 40.5% after fine-tuning.
@@
EvoEval includes 828 programming problems created by prompting GPT-4 to evolve original HumanEval tasks across 5 semantic-altering and 2 semantic-preserving benchmarks, each of which has 100 problems. We conduct experiments on the Tool_Use, Subtle, and Creative benchmarks to evaluate the performance of Qwen2.5-Coder-7B-Instruct and deepseek-coder-6.7b-instruct fine-tuned with . As shown in Table 4, both models demonstrate significant improvements in code efficiency after fine-tuning with . For the Tool_Use benchmark, the average execution time (ET) for correct code generated by Qwen2.5-Coder-7B-Instruct decreases by 42.9%, and the total memory usage (TMU) decreases by 41.6% after fine-tuning. Similarly, deepseek-coder-6.7b-instruct achieves a 41.5% reduction in ET and a 15.9% reduction in TMU. In the Subtle benchmark, after fine-tuning, Qwen2.5-Coder-7B-Instruct and deepseek-coder-6.7b-instruct exhibit 17.5% and 33.3% reductions in ET, respectively, after fine-tuning. The normalized total memory usage (NTMU) also decreases by 19.9% and 34.5% for the two models. Moreover, the pass@1 rate improves significantly, with Qwen2.5-Coder-7B-Instruct increasing from 55.0% to 72.0% and deepseek-coder-6.7b-instruct increasing from 56.0% to 69.0%. For the Creative benchmark, Qwen2.5-coder-7B-instruct and deepseek-coder-6.7b-instruct achieve 27.5% and 38.1% reductions in ET, respectively, after fine-tuning. The NTMU also decreases by 23.8% and 39.5% for the two models. The pass@1 rate remains the same for Qwen2.5-Coder-7B-Instruct at 44.0% but improves from 31.0% to 41.0% for deepseek-coder-6.7b-instruct.
@@ 4.2 Data Science Programming
DS-1000 is a data science benchmark consisting of 1000 realistic challenges across 7 popular Python data science libraries. We evaluate the efficiency and pass@1 of LLM-generated code for the DS-1000 tasks using Qwen2.5-Coder-7B-Instruct and deepseek-coder-6.7b-instruct, both in their original form and fine-tuned with . As shown in Table 5, fine-tuning with leads to modest improvements in code efficiency for both models. For Qwen2.5-Coder-7B-Instruct, the average execution time (ET) decreases by 0.8%, and the total memory usage (TMU) decreases by 0.9% after fine-tuning, while the pass@1 rate improves significantly from 17.00% to 35.70%. For deepseek-coder-6.7b-instruct, fine-tuning results in a 4.5% reduction in ET and a 5.6% reduction in TMU, with the memory usage (MU) and normalized memory usage (NMU) also decreasing by 1.9% and 2.0%, respectively, although the pass@1 rate remains the same at 37.50%. While the improvements in code efficiency for the DS-1000 benchmark are less pronounced compared to other benchmarks, the results still demonstrate that fine-tuning with can enhance the efficiency of LLM-generated code for data science tasks.
@@ 4.3 Different Programming Language
In addition to evaluating the efficiency and pass@1 of LLM-generated code for Python tasks, we also conduct experiments on the HumanEval-X dataset, where we focus on the C++ and Java subsets, to measure the performance of fine-tuned LLMs in a different-language setting. As shown in Table 6, both Qwen2.5-Coder-7B-Instruct and deepseek-coder-6.7b-instruct demonstrate significant improvements in code efficiency and pass@1 rates after fine-tuning with for C++ tasks. Qwen2.5-Coder-7B-Instruct achieves a 37.1% reduction in average execution time (ET), a 42.6% reduction in memory usage (MU), and a 53.8% reduction in total memory usage (TMU), with the pass@1 rate improving from 7.32% to 48.17%. Similarly, deepseek-coder-6.7b-instruct exhibits a 45.2% reduction in ET, a 50.0% reduction in MU, and a 40.0% reduction in TMU, with the pass@1 rate increasing from 14.63% to 40.24%. For Java tasks, deepseek-coder-6.7b-instruct achieves a 36.5% reduction in ET, a 50.5% reduction in MU, and a 30.4% reduction in TMU after fine-tuning, with the pass@1 rate improving from 14.63% to 57.93%, matching the performance of the fine-tuned Qwen2.5-Coder-7B-Instruct. These results demonstrate that fine-tuning with can significantly enhance the efficiency and correctness of LLM-generated code across multiple programming languages.